# A search for Galactic post-AGB stars in Gaia DR3††thanks: Tables A6 and A7 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A

### v1

@@ 3.4 Galactic heights and membership to the Halo
Gaia DR3 precise distances allowed us to calculate the galactic distribution of our final sample. Figure 8 depicts the galactic height as a function of the galactic longitude. It also shows the commonly adopted limits for the main structures in the Milky Way: thin disk, thick disk, and halo.
This distribution allows us to tentatively assign the 118 objects either to the disk 83 objects with kpc) or to the halo (35). Suspected halo stars are flagged with ’H’ in Tables A2, A3, and A5. Although this classification is adopted to compare their position in the HR diagram with evolutionary tracks suited for each of the populations, we are well aware that it will benefit from a spectroscopic confirmation.
@@ 3.5 Classification: identification of post-AGB stars and other objects
Once luminosity values and their uncertainties are known, we can apply the luminosity thresholds discussed before. We obtain that 69 objects can be classified as post-AGB stars bona fide candidates, 46 objects can not be confirmed as post-AGB stars because we derive a luminosity lower than 2500 for them, and 3 objects above the high 35,000 luminosity threshold are classified as supergiant stars.
@@
From the sample of 46 stars with luminosities lower than 2500 , 5 are found to be YSOs in Molecular Clouds (see further in this section), 3 are suspected or confirmed to be Horizontal Branch (HB) stars, and 38 remain as unclassified. From this last group, 9 objects with luminosity lower than 100 are tentatively classified as possible YSOs as will be discussed later. The properties of the objects in the main categories are listed in Table A.2 (post-AGB stars), Table A.3 (unconfirmed post-AGB candidates), Table A.4 (YSOs candidates), and Table A.5 (supergiants and HB stars).
@@
To illustrate these results, we have depicted all these objects in the HR diagram (Fig. 9) together with the evolutionary tracks for post-AGB stars and PNe central stars from Miller Bertolami (2016). To derive the masses we used the tracks with metallicity of , for stars with kpc. expected to be in the disc (upper panel), while for those with kpc we used the tracks. Using the luminosity threshold for post-AGB stars discussed before, the diagram allowed us to disclose bona fide post-AGB candidates from those that are not.
@@
We would like to stress the fact that we are applying a quite restrictive selection threshold to sort out bona fide post-AGB candidates. It could certainly be the case that some other objects, located close to our luminosity threshold, could also be post-AGB but, given the uncertainties, do not fulfil our selection criteria, that we named as possible post-AGB stars.
@@
It is worth noticing here that 12 of the objects that we classify as post-AGBs have been identified as possible, or confirmed, binary stars by Kluska et al. (2019). We checked the bibliography for other binary references, and found 5 additional binary stars. These stars together with two possible YSO binary stars, also according to Kluska et al. (2019), one supergiant star and one post-AGB unconfirmed candidate are identified with a flag in Table LABEL:table:candidates. It is worth noticing as well, that our SED classification is disc-type for all but one of those binary objects.
@@
Figure 9 also shows two types of YSOs. All those objects with are suspected to be YSOs, but they could also be other types of evolved stars such as HB stars. There are also examples of well-known YSOs among objects with higher luminosities. The physical nature as YSOs of those luminous objects is more difficult to discern because their luminosities and temperatures overlap with those of post-AGB stars, and what we can do is resort to studying their locations in the Galaxy. If their positions and distances match those of star-forming regions, they can quite safely be classified as young objects.
@@
We used the molecular clouds catalogue by Zucker et al. (2020), which gives coordinates and distances to a large number of those regions, and found that 5 of our objects are located within these clouds. Such objects are labeled as YSO in MoC (Molecular Clouds) in Figure 9 and Table A.4.
@@
From those, 3 objects have luminosities above and 2 below that limit. The classification as YSOs of the remaining 8 objects with luminosities below is tentative, and consequently, we label them as possible YSOs in Fig. 9.
@@
Figure 9 also shows 3 objects that we found to be Horizontal Branch stars: SDS2012 Ter8 38 is a blue Horizontal Branch star in the globular cluster Terzan 8, [SDS2012] NGC 6402 160 in NGC 6402 and BPS BS 16479-0009 is a field Horizontal Branch star candidate according to Beers et al. (1996). Finally, we found three objects above the luminosity upper limit expected for post-AGB stars that we classify as supergiant stars. We shall comment further on them in section 5.
@@ 3.6 Comparison with other classifications in the literature
We can compare our classification with that recently obtained by Aoki et al. (2022) for the 7 objects in common in both samples. 4 post-AGBs are identically classified by both of us, while 2 of our unconfirmed post-AGB stars were catalogued as post-AGB or as cool post-AGB by them and one of our possible YSOs was catalogued as a hot subdwarf by those authors.
When comparing our results with those in Kamath et al. (2022) we found 11 objects in common (those with good astrometric quality in that work, RUWE¡1.4). We can confirm their nature as bona fide post-AGB candidates for 10 of them while HD 107360 remains slightly underluminous for the post-AGB threshold. In general, the temperatures and luminosity values given by Kamath et al. (2022) are in agreement with our derived values, as it is illustrated in Table 1.
@@
In Section 3.3 we compared our results with those obtained by Oudmaijer et al. (2022), as shown in Figures 5 and 7. For the Luminosity range expected for post-AGB stars, temperatures and luminosities agree with a dispersion that can be explained by methodological differences, as already discussed in that section. From 59 objects in common with Oudmaijer et al. (2022) sample, we have classified as post-AGB 44 of them, while we have catalogued 8 of them as unconfirmed candidates, 3 as possible YSOs, 2 as YSOs in molecular clouds and another 2 as supergiants.
@@
Finally, we can summarise our results. Starting from the lists of post-AGB objects known or proposed as such in the literature, we have made a careful selection based on the quality of the astrometry in Gaia DR3 of these sources. We used updated dust extinction maps, 3D when available, for a more accurate derivation of luminosities. As a consequence, we were able to classify some of them as bona-fide post-AGB stars (69), supergiants (3), HB stars (3), YSOs in molecular clouds (5) and possible YSOs (9), while 29 objects remain as unconfirmed post-AGB candidates. In the following section, we describe the evolutionary properties of our sample of 69 post-AGB stars.
4 Sample of post-AGB stars
@@
We focussed now on the sample of 69 objects whose luminosities allow us to confirm their evolutionary state as bona fide post-AGB candidates. By interpolating between the novel evolutionary models by Miller Bertolami (2016) for post-AGB stars, we estimate their progenitor mass (in the MS) and their evolutionary age in the post-AGB phase. Miller Bertolami (2016) provides tracks for metallicity values 0.01, 0.02, 0.001, and 0.0001. As we already mentioned, we used the tracks for as representative of the disc population, and for those objects belonging to the halo we used tracks.
@@
Note that the objects classified as unconfirmed post-AGB candidates are located below the 1 track (0.9 for halo stars). It is assumed that these objects, if they are single-evolved stars, can only have initial masses slightly below 1 . Conversely, they could have their origin in binary evolution.
@@
The progenitor mass distribution we obtained for the post-AGB stars in our sample is displayed in Fig. 10. About half of our stars (35) have masses below 1.5 in the MS, while only 5 of them have a mass higher than 3.5 . Although our study is limited in the number of objects and possibly comes from a biased selection, the resulting masses match the expected distribution. Parent stars with masses above 3.5 have too short lifetimes in the post-AGB phase (the crossing times for post-AGB and PNe phases in Miller Bertolami tracks are less than 400 yr) to populate this region. The mean value of the progenitor masses for the post-AGB sample is:
@@
This mean mass value agrees with the value of obtained in González-Santamaría et al. (2021) for stars in the next evolutionary phase, as central stars of planetary nebulae (CSPNe).
The mean mass value for the post-AGB stars final sample resulted:
@@
Regarding evolutionary ages, we obtained the distribution shown in Fig. 11. Around 36% of the stars have ages smaller than 2000 yrs. Moreover, almost all the objects in the sample show evolutionary ages below 10,000 yrs, with only 6 stars over that age. We have obtained the following evolutionary age mean value of the sample:
@@
To be reminded is that in Miller Bertolami models, the beginning of the post-AGB phase is taken when, due to stellar winds, the mass of the external layer of the star drops below 1% of the star mass.
@@ 5 Special objects
It deserves special mention that our study allows confirming the nature as post-AGB objects of three objects, lacking a previous firm classification in the catalogues, neither in the Torun catalogue as Likely, nor in Simbad or in the Suárez et al. catalogues as confirmed post-AGB.
CD-30 15464 – This object is catalogued as Possible post-AGB in the Torun catalogue, but as simple star in Simbad and it has no references in Suárez et al. (2006) catalogue. It is reported as a B1 spectral type star, while its effective temperature is quite high (22,000 K), still within the limits of a post-AGB star. It is located very far from the Sun, at a distance of 9.96 kpc, at the galactic centre direction () and slightly below the galactic disk (). Despite this location, the interstellar extinction is not very high in the direction of this star (), and its image777http://cdsportal.u-strasbg.fr/?target=CD-30%2015464 in PanSTARRS colours shows evidence of a circumstellar envelope.
HD 53300 – It is catalogued as A2 type star and it is located in the galactic disk (). Instead, Rao et al. (2012) derived an effective temperature of 7250 K for it, which corresponds to a F1 type star. This star is mentioned as a candidate post-AGB in the Simbad database and it has no references in the Torun or Suárez et al. (2006) catalogues, although Rao et al. (2012) classifies it as post-AGB star, while in Bhatt & Manoj (2000) it is identified as a Vega-like star.
HD 214539 – This star is catalogued as Possible post-AGB in Torun and as simple Star in Simbad, with a B8/9 spectral type. Kodaira & Philip (1984) obtained a temperature of 10,000 K and a gravity of log[g]=2 for this object. It is located at a distance of 1.39 kpc. By a visual inspection888http://cdsportal.u-strasbg.fr/?target=HD%20214539, also in this case, a circumstellar envelope typical of post-AGB stars can be observed.
@@
BD-02 4931 – This object is catalogued as post-AGB candidate star in the Simbad database, while in Parthasarathy et al. (2000a) it is classified as a B1 type giant. Assuming a temperature of 16,000 K from its spectral type (the same value as the one obtained by fitting its SED by Gielen et al. 2011) we obtain a luminosity of , higher than the predictions for post-AGB stars.
HD 179821 – This star is catalogued as Likely post-AGB in the Torun catalogue and as a post-AGB star in the Simbad database and in Suárez et al. catalogue. However, we have obtained a high luminosity of (very similar to that obtained by Wood et al. (1983), ) and low surface gravity of for this object, that leads us to classify it as a supergiant star. Moreover, in Şahin et al. (2016), they catalogued it as a likely massive post-red supergiant star.
V* V1027 Cyg – This object is also classified as Likely post-AGB in the Torun catalogue and as post-AGB in Simbad database. But, in this case, we obtained stellar parameters typical of supergiant stars, such as a high luminosity of and a very low surface gravity of . Furthermore, in the Winfrey et al. (1994) catalogue, they classified it as a G7 supergiant star, while it is catalogued as a G8-K3 type supergiant in a more recent study by Arkhipova et al. (2016).
@@ 6 Conclusions
From a sample of 118 post-AGB stars candidates selected from the literature and having filtered out those with Gaia DR3 not accurate enough astrometry (and hence unreliable distances), we have estimated their luminosity values which allow us to classify them as bona fide post-AGB candidates (69) or as objects in a different evolutionary phase, such as YSOs (5), possible YSOs (9), supergiant stars (3) and HB stars (3), while 29 stars remain as unconfirmed post-AGB candidates.
@@
Using the spanish VOSA service, we have fitted the SEDs of each star and simultaneously obtained its effective temperature and luminosity. This allows us to plot the post-AGB candidates in an HR diagram, and by using Miller Bertolami (2016) evolutionary tracks in the post-AGB phase, to derive their masses and ages. We found that our 69-object sample includes mainly stars with progenitor masses between 1 and 2.5 values that agree with the type of post-AGB stars that statistically could be found in a small sample like ours. The mass mean value of the sample is also in agreement with that expected for stars in the planetary nebula phase (the next evolutionary phase), according to González-Santamaría et al. (2021).
@@
Our study allows us to confirm the nature as post-AGB stars of several objects previously unconfirmed as such. It is important to note that although the Gaia DR3 catalogue contains statistically valuable information for an enormous number of stars, many of the available parallaxes contain errors too large to allow a correct estimation of the absolute magnitude or luminosity of the stars. We chose to perform rather restrictive filtering of the astrometric quality of Gaia measurements (parallax and distance errors, RUWE value) as well as of the distance inferred from the parallax measurement by a Bayesian model of galactic stellar distributions.
@@
Behind this filtering is the idea of working with a small set of objects, candidates to be in the post-AGB phase, all of them with precise distances in Gaia, which therefore will mostly be individual objects for which a luminosity can be calculated with great confidence. For this reason, our methodology was based on selecting only a subset of stars that are firm candidates to be in the post-AGB phase and exhibit Gaia DR3 parallaxes (and inferred distances) with errors below 30%.
@@
Also to be noticed is that we worked only with those candidates with interstellar extinction values from Stassun 3D dust maps, which allowed us to initially constrain interstellar extinction. Total extinction, including both interstellar and circumstellar extinction, was derived simultaneously with the luminosity value from the SED fitting procedure. Effective temperatures were taken from the literature, using spectroscopically determined temperatures when they were available.
@@
We discuss in detail our results in comparison with other similar studies about post-AGBs, carried out recently. About 25% of the stars in our sample resulted underluminous to be confirmed as post-AGB stars under our working hypotheses, while 12% are found to be possible YSOs, either located in Molecular Clouds (5) or candidates (9). Other objects such as HB stars (3) or supergiant stars (3) were also included in the original compilation.
@@
Regarding binarity, although our initial filtering would rule out most of the binary objects, we found that 18 objects out of our sample of 69 post-AGB have SEDs that can be classified as disk-type and, therefore, might correspond to binary objects. Searching the literature for binarity, we found that 17 objects among them have been identified as such either by Kluska et al. 2022 (12 objects) or by other authors (see table A.1 for references). This brings us to some 18 possible or confirmed binaries within our sample, about 26%.
@@
We think our results provide an interesting framework to pursue further insight into the study of the post-AGB phase. In particular, given our well characterised sample of 69 objects, a way is open to complement the study on the unconfirmed 29) candidates proposed. A follow-up analysis of their properties, including spectroscopy when possible, would be desirable.

### v2

@@ 3.4 Galactic heights and membership to the halo
The precise Gaia DR3 distances allowed us to calculate the Galactic distribution of our final sample. Figure 8 depicts the Galactic height as a function of the Galactic longitude. It also shows the commonly adopted limits for the main structures in the Milky Way: the thin disc, thick disc, and the halo.
This distribution allowed us to tentatively assign the 118 objects to either the 83 disc objects with kpc) or to the halo (35). Suspected halo stars are flagged with ’H’ in Tables A2, A3, and A5. Although this classification was adopted to compare their position in the HR diagram with evolutionary tracks suited for each of the populations, we are well aware that it will benefit from a spectroscopic confirmation.
@@ 3.5 Classification: Identification of post-AGB stars and other objects
After the luminosity values and their uncertainties were known, we applied the luminosity thresholds discussed before. We obtained that 69 objects can be classified as post-AGB star bona fide candidates, 46 objects cannot be confirmed as post-AGB stars because we derived a luminosity lower than 2500 for them, and 3 objects above the high 35,000 luminosity threshold are classified as supergiant stars.
@@
In the sample of 46 stars with luminosities lower than 2500 , 5 are found to be YSOs in molecular clouds (see below), 3 are suspected or confirmed to be Horizontal Branch (HB) stars, and 38 remain unclassified. In this last group, 9 objects with luminosity lower than 100 are tentatively classified as possible YSOs, as we discuss below. The properties of the objects in the main categories are listed in Table A.2 (post-AGB stars), Table A.3 (unconfirmed post-AGB candidates), Table A.4 (YSOs candidates), and Table A.5 (supergiants and HB stars).
@@
To illustrate these results, we depicted all these objects in the HR diagram (Fig. 9) together with the evolutionary tracks for post-AGB stars and PNe central stars from Miller Bertolami (2016). To derive the masses, we used the tracks with a metallicity of for stars with kpc that are expected to be in the disc (upper panel), while for those with kpc, we used the tracks. With the luminosity threshold for post-AGB stars discussed before, the diagram allowed us to disclose bona fide post-AGB candidates from those that are not.
@@
We would like to stress the fact that we applied a quite restrictive selection threshold to sort out bona fide post-AGB candidates. Some other objects that are located close to our luminosity threshold might also be post-AGB, but given the uncertainties, they do not fulfil our selection criteria for possible post-AGB stars.
@@
Twelve of the objects that we classified as post-AGBs have been identified as possible or confirmed binary stars by Kluska et al. (2019). We checked the bibliography for other binary references and found 5 additional binary stars. We identify these stars, together with two possible YSO binary stars, also according to Kluska et al. (2019), one supergiant star, and one post-AGB unconfirmed candidate with a flag in Table LABEL:table:candidates. Moreover, our SED classification is disc-type for all but one of these binary objects.
@@
Figure 9 also shows two types of YSOs. All objects with are suspected to be YSOs, but they might also be other types of evolved stars, such as HB stars. There are also examples of well-known YSOs among the objects with higher luminosities. The physical nature of YSOs of these luminous objects is more difficult to discern because their luminosities and temperatures overlap with those of post-AGB stars. We studied their locations in the Galaxy. If their positions and distances matched those of star-forming regions, they were quite safely classified as young objects.
@@
We used the molecular cloud catalogue by Zucker et al. (2020), which gives coordinates and distances to a large number of these regions. We found that five of our objects are located within these clouds. These objects are labelled ”YSO in MoC” (molecular clouds) in Figure 9 and Table A.4.
@@
The luminosities of three of these objects are above , and the luminosities of two others are below this limit. The classification as a YSO of the remaining eight objects with luminosities below is tentative. We therefore label them possible YSOs in Fig. 9.
@@
Figure 9 also shows three objects that we found to be Horizontal Branch stars: SDS2012 Ter8 38 is a blue Horizontal Branch star in the globular cluster Terzan 8, [SDS2012] NGC 6402 160 in NGC 6402 and BPS BS 16479-0009 is a field Horizontal Branch star candidate according to Beers et al. (1996). Finally, we found three objects above the upper luminosity limit expected for post-AGB stars that we classify as supergiant stars. We comment further on them in section 5.
@@ 3.6 Comparison with other classifications in the literature
We can compare our classification with the classification recently obtained by Aoki et al. (2022) for the seven objects in common in both samples. Four post-AGBs are identically classified by both of us, while two of our unconfirmed post-AGB stars were catalogued as post-AGB or as cool post-AGB by them, and one of our possible YSOs was catalogued as a hot subdwarf by these authors.
When comparing our results with those in Kamath et al. (2022), we found 11 objects in common (those with good astrometric quality in that work, RUWE¡1.4). We can confirm a nature as bona fide post-AGB candidates for 10 of these, while HD 107360 remains slightly underluminous for the post-AGB threshold. In general, the temperatures and luminosity values given by Kamath et al. (2022) agree with our derived values, as ,illustrated in Table 1.
@@
In Section 3.3 we compared our results with those obtained by Oudmaijer et al. (2022), as shown in Figures 5 and 7. For the luminosity range expected for post-AGB stars, the temperatures and luminosities agree with a dispersion that can be explained by methodological differences, as already discussed in that section. Forty-four of the 59 objects in common with Oudmaijer et al. (2022) sample were classified as post-AGB 44, and we catalogued 8 of them as unconfirmed candidates, 3 as possible YSOs, 2 as YSOs in molecular clouds, and another 2 as supergiants.
@@
Finally, we summarise our results. Starting from the lists of post-AGB objects known or proposed as such in the literature, we selected stars based on the quality of the astrometry in Gaia DR3 of these sources. We used updated dust-extinction maps, 3D when available, to derive more accurate luminosities. As a consequence, we were able to classify some of them as bona fide post-AGB stars (69), supergiants (3), HB stars (3), YSOs in molecular clouds (5), and possible YSOs (9), while 29 objects remain unconfirmed post-AGB candidates. In the following section, we describe the evolutionary properties of our sample of 69 post-AGB stars.
4 Sample of post-asymptotic giant branch stars
@@
We focus now on the sample of 69 objects whose luminosities allowed us to confirm their evolutionary state as bona fide post-AGB candidates. By interpolating between the novel evolutionary models by Miller Bertolami (2016) for post-AGB stars, we estimated their progenitor mass (in the MS) and their evolutionary age in the post-AGB phase. Miller Bertolami (2016) provided tracks for metallicity values 0.01, 0.02, 0.001, and 0.0001. We used the tracks for as representative of the disc population, and for objects belonging to the halo, we used tracks.
@@
The objects classified as unconfirmed post-AGB candidates are located below the 1 track (0.9 for halo stars). It is assumed that the initial masses of these objects, if they are single-evolved stars, can only be slightly below 1 . Conversely, they could have their origin in binary evolution.
@@
The progenitor mass distribution we obtained for the post-AGB stars in our sample is displayed in Fig. 10. The masses of about half of our stars (35) are below 1.5 in the MS, and the masses of only 5 of them are higher than 3.5 . Although our study is limited in the number of objects and possibly comes from a biased selection, the resulting masses match the expected distribution. The lifetimes of parent stars with masses above 3.5 are too short in the post-AGB phase (the crossing times for post-AGB and PNe phases in Miller Bertolami tracks are shorter than 400 yr) for them to populate this region. The mean value of the progenitor masses for the post-AGB sample is
@@
This mean mass value agrees with the value of obtained in González-Santamaría et al. (2021) for stars in the next evolutionary phase, as central stars of planetary nebulae.
The mean mass value for the post-AGB stars final sample is
@@
We obtained the evolutionary age distribution shown in Fig. 11. The ages of about 36% of the stars are younger than 2000 yr. Moreover, almost all the objects in the sample show evolutionary ages below 10,000 yr, and only 6 stars are older. We obtained the following mean evolutionary age of the sample:
@@
In Miller Bertolami models, the beginning of the post-AGB phase is taken when the mass of the external layer of the star drops below 1% of the star mass as a result of stellar winds.
@@ 5 Special objects
Our study allows us to confirm the nature of three objects as post-AGB stars. They lacked a previous firm classification in the catalogues and were neither listed in the Torun catalogue as likely, nor in Simbad or in the Suárez et al. catalogues as confirmed post-AGB.
CD-30 15464. This object is catalogued as possible post-AGB in the Torun catalogue but as simple star in Simbad, and it has no references in Suárez et al. (2006) catalogue. It is reported as a star of spectral type B1, while its effective temperature is quite high (22,000 K), still within the limits of a post-AGB star. It is located very far from the Sun, at a distance of 9.96 kpc, in the direction of the Galactic centre () and slightly below the Galactic disc (). Despite this location, the interstellar extinction is low in the direction of this star (), and its image131313http://cdsportal.u-strasbg.fr/?target=CD-30%2015464 in PanSTARRS colours shows evidence of a circumstellar envelope.
HD 53300 is catalogued as an A2 type star and is located in the Galactic disc (). Instead, Rao et al. (2012) derived an effective temperature of 7250 K for it, which corresponds to an F1 type star. This star is mentioned as a candidate post-AGB in the Simbad database and lacks references in the Torun or Suárez et al. (2006) catalogues, although Rao et al. (2012) classified it as post-AGB star, while it was classified in Bhatt & Manoj (2000) as a Vega-like star.
HD 214539.This star is catalogued as possible post-AGB in Torun and as simple star in Simbad, with a spectral type B8/9. Kodaira & Philip (1984) obtained a temperature of 10,000 K and a gravity of log[g]=2 for this object. It is located at a distance of 1.39 kpc. By visual inspection141414http://cdsportal.u-strasbg.fr/?target=HD%20214539, a circumstellar envelope typical of post-AGB stars can also be observed.
@@
BD-02 4931. This object is catalogued as a post-AGB candidate star in the Simbad database, while in Parthasarathy et al. (2000a), it is classified as a B1 type giant. Assuming a temperature of 16,000 K from its spectral type (the same value as was obtained by fitting its SED by Gielen et al. 2011), we obtained a luminosity of , which is higher than the predictions for post-AGB stars.
HD 179821. This star is catalogued as likely post-AGB in the Torun catalogue and as a post-AGB star in the Simbad database and in the Suárez et al. catalogue. However, we obtained a high luminosity of (very similar to that obtained by Wood et al. (1983), ) and low surface gravity of for this object, which led us to classify it as a supergiant star. Moreover, Şahin et al. (2016) catalogued it as a likely massive post-red supergiant star.
V* V1027 Cyg. This object is also classified as likely post-AGB in the Torun catalogue and as post-AGB in Simbad database. In this case, however, we obtained stellar parameters typical of supergiant stars, such as a high luminosity of and a very low surface gravity of . Furthermore, in the Winfrey et al. (1994) catalogue, they classified it as a G7 supergiant star, but it was catalogued as a G8-K3 type supergiant in a more recent study by Arkhipova et al. (2016).
@@ 6 Conclusions
Based on a sample of 118 post-AGB star candidates selected from the literature and after filtering out stars for which Gaia DR3 astrometry was not accurate enough (and for which the distances were therefore unreliable), we have estimated their luminosity values, which allowed us to classify them as bona fide post-AGB candidates (69) or as objects in a different evolutionary phase, such as YSOs (5), possible YSOs (9), supergiant stars (3), and HB stars (3). Twenty-nine stars remain unconfirmed post-AGB candidates.
@@
Using the Spanish VOSA service, we fitted the SEDs of each star and simultaneously obtained its effective temperature and luminosity. This allowed us to plot the post-AGB candidates in an HR diagram, and by using the Miller Bertolami (2016) evolutionary tracks in the post-AGB phase, we derived their masses and ages. We found that our 69-object sample includes mainly stars with progenitor masses between 1 and 2.5 , which agree with the type of post-AGB stars that statistically could be found in a small sample like ours. The mass mean value of the sample also agrees with that expected for stars in the planetary nebula phase (the next evolutionary phase) according to González-Santamaría et al. (2021).
@@
Our study allowed us to confirm the nature of several objects as post-AGB stars that were previously not confirmed as such. It is important to note that although the Gaia DR3 catalogue contains statistically valuable information for an enormous number of stars, many of the available parallaxes contain errors that are too large to allow a correct estimation of the absolute magnitude or luminosity of the stars. We chose to perform a rather restrictive filtering of the astrometric quality of Gaia measurements (parallax and distance errors, RUWE value) as well as of the distance inferred from the parallax measurement by a Bayesian model of Galactic stellar distributions.
@@
This filtering was based on the idea of working with a small set of objects, candidates for the post-AGB phase, all of which have precise distances in Gaia, which therefore mostly are individual objects for which a luminosity can be calculated with great confidence. For this reason, our method was based on selecting only a subset of stars that are firm candidates to be in the post-AGB phase and have Gaia DR3 parallaxes (and inferred distances) with errors below 30%.
@@
We also only worked with candidates with interstellar extinction values from Stassun 3D dust maps, which allowed us to initially constrain interstellar extinction. The total extinction, including both interstellar and circumstellar extinction, was derived simultaneously with the luminosity value from the SED-fitting procedure. The effective temperatures were taken from the literature, using spectroscopically determined temperatures when they were available.
@@
We discussed our results in comparison with other similar studies of post-AGBs that were carried out recently. About 25% of the stars in our sample are too underluminous to be confirmed as post-AGB stars under our working hypotheses, while 12% are found to be possible YSOs, are either located in Molecular Clouds (5) or are candidates (9). Other objects, such as HB stars (3) or supergiant stars (3), were also included in the original compilation.
@@
Although our initial filtering would rule out most of the binary objects, we found that the SEDs of 18 objects out of our sample of 69 post-AGB can be classified as disc-type. They might therefore be binary objects. Searching the literature for binarity, we found that 17 of these objects were identified as such either by Kluska et al. 2022 (12 objects) or by other authors (see table A.1 for references). This means that our sample contains 18 possible or confirmed binaries, which is about 26%.
@@
Our results provide an interesting framework for further insight into the post-AGB phase. In particular, our well-characterised sample of 69 objects opens the way to complementing the study on the unconfirmed 29 candidates. A follow-up analysis of their properties, including spectroscopy when possible, would be desirable.